A rectangle with an area of 125cm2 has sides in the ratio 4:5. What is the perimeter of the rectangle?

Answer by JulietG(1812)

The area is length x width.
L * W = 125
What are the factors of 125?
125, 1
5, 25
Neither of those has a ratio of 4:5. It's not a whole number.
.
It can still be solved.
Let the sides be 4x and 5x. (ratio 4:5)
125 = 4x * 5x
125 = 20x^2
x^2 = 125/20
x^2 = 6.25
Take the square root.
x = sqaure root of 6.25
x = 2.5
4x = 4 * 2.5 = 10
5x = 5 * 2.5 = 12.5
The perimeter is a fence. You'll need twice the length (12.5) and twice the width (10)
(12.5*2)+ (10*2)
25 + 20 =
45